OpenStudy (anonymous):

Someone please help me solve this!!! f^-1(3) when f(x)=2x+3/5

OpenStudy (solomonzelman):

you function is: I can just replace f(x) with a 3.

OpenStudy (solomonzelman):

plug in 3 for f(x) (not f(3), but set f(x)=3 )

OpenStudy (anonymous):

Okay, does this mean that it would be 2(3)+3/5??

OpenStudy (solomonzelman):

no, it means that it is 3=2x+3/5

OpenStudy (anonymous):

OHHHHHHH okay so now I do what?

OpenStudy (solomonzelman):

soilve for x.

OpenStudy (anonymous):

So solve 2x+3 first? then divide it by 5? is that right?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

is it y=(2x+3)/5 , it was that?

OpenStudy (anonymous):

f(x)=2x+3/5

OpenStudy (solomonzelman):

so 2x is NOT part of the fraction ?

OpenStudy (isaiah.feynman):

Find the inverse of f(x) then evaluate it at 3.

OpenStudy (solomonzelman):

No

OpenStudy (solomonzelman):

it is easier to plug in 3 instead of y. Isn't it, Isaiah?

OpenStudy (anonymous):

F(x)= 2x+3 ---------- 5

OpenStudy (solomonzelman):

Oh, yes, so we are at: 3=(2x+3) / 5 multiply both sides by 5.

OpenStudy (solomonzelman):

what do you get when you multiply both sides by 5, can you tell me?

OpenStudy (anonymous):

So would it be.... 5x=2x+3???

OpenStudy (isaiah.feynman):

@SolomonZelman I only know the traditional method.

OpenStudy (solomonzelman):

If you need the f^(-1)(x) then do the traditional, but for this, I think you see why plugging in 3 for y, is still same as f^(-1)(3).

OpenStudy (anonymous):

Okay, am i right with what i said for multiplying by 5?

OpenStudy (solomonzelman):

@nicjohns 3=(2x+3)/5 3* 5 = 5* (2x+3)/5 15=2x+3

OpenStudy (anonymous):

Okay and that is my final answer or i solve for x by one step equation?

OpenStudy (anonymous):

So subtract 3 from both sides.... 12=2x. Divide by 2 on both sides and equal x=6????

OpenStudy (solomonzelman):

you have to solve for x. isolate the x first by subtracting 3 from both sides, and then by dividing both sides by 2.

OpenStudy (anonymous):

I just did that, and put it! x=6!!!!

OpenStudy (solomonzelman):

yes, that is right!

OpenStudy (anonymous):

Thank you so much!! I fi open another question so I can give you another medal, can you help me with one more???

OpenStudy (solomonzelman):

I will see if I get more time to help you. yw:)